OpenStudy (anonymous):

Having a problem with algebra (just a simple question that I should know the answer to, but I'm just having one of those days). Could somebody help?

OpenStudy (anonymous):

The answer I marked is what I think is correct.

zepdrix (zepdrix):

Hmm I had to Google it, I'm not really familiar with the term "radical equation" either. It looks like it just means anything involving a `radical` on the x. Examples:\[\Large\rm \sqrt x+2=7\]\[\Large\rm 4\sqrt x=24\]

OpenStudy (anonymous):

So then it would be d instead of a?

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

look at where the x is...

zepdrix (zepdrix):

Yes, d :)